Determine the bandwidth for a communication channel that has an SNR of 35 dB and capable of transmitting the data rate at 37.2 Kbps.
Answer #1

Bandwidth, SNR and data rate is related in below formula:

Shannon Formula: C = B * log2(1+ SNR)

where C data transfer rate, in bits/s

B is the bandwidth of the line, in bits/s

S is the average signal power and

N is the average noise power.

The signal-to-noise ratio (SNR) is expressed in decibels (dB) question.

Given,

Data transfer rate, C = 37.2 Kbps => 37200 bps

SNR = 35 dB

Bandwidth, B = ?

According to formula,

B = C / log2(1+ SNR)

Putting values,

B = 37200 * log2( 1 + 35 )

B = 37200 * log2( 36 )

B = 37200 * 5.1699

B = 192320.28 Hz

B = 192.3203 KHz

Required bandwidth is 192.3203 KHz.